How many different ways can 3 bus drivers be selected from 11 bus drivers for a public transport service in a city


1
Expert's answer
2021-06-08T03:40:14-0400

N=11

n=3

Number of ways "= \frac{N!}{n!(N-n)}"

"= \frac{11!}{3!(11-3)!} \\

= \frac{990}{6} \\

=165"